Evolution:

Context-4/4 pts Question 1 A sample of columbine plants from a natural population are genotyped for a polymorphism in the R gene, whi0/4 pts Incorrect Question 2 In the sample described in the previous question, what would be the expected frequency of the R1
Answer #1

As per HWE. p + 9 = 1 p^2 + 2pq + q^2 = 1 p^2 = genotypic frequency of homozygous dominant genotype 2pq = genotypic frequency

Total individuals = 28 + 64 + 108 = 200

GF of R1R1 = 0.3 × 0.3 = 0.09

AF of R1 = √0.14 = 0.3745 = 0.3 approx.

Please rate.

As per HWE. p + 9 = 1 p^2 + 2pq + q^2 = 1 p^2 = genotypic frequency of homozygous dominant genotype 2pq = genotypic frequency of heterozygotes q^2 = genotypic frequency of homozygous recessive genotype p = allelic frequency of dominant allele q = allelic frequency of recessive allele Number of individuals with a particular genotype = total number of individuals x genotypic frequency of that genotype